A person wants to make a one ounce ring composed of Gold and Silver. One ounce of gold cost $600 and one ounce of silver cost $50. How much of each metal should be used if the cost is $200
Answer by stanbon(75887)

A person wants to make a one ounce ring composed of Gold and Silver. One ounce of gold cost $600 and one ounce of silver cost $50. How much of each metal should be used if the cost is $200
----------------
Weight Eq.::: g + s = 1 ounce
Value Eq. :::600g + 50s = 200
--------------------------------------
Multiply thru the Weight Eq by 50 to get:
50g + 50s = 50
---
Subtract that from the Value Eq. and solve for "g":
550g = 150
g = 3/11 ounce (amt of gold needed in the mixture)
----
Since g + s = 1, s = 8/11 ounce (amt. of silver needed in the mixture)
==========================================================================
Cheers,
Stan H.